OpenStudy (anonymous):

fluid mixing question, i calculated the rate in - rate out to be 500-(y(t)/(40+t)). it's now asking me to solve the differential equation given y(0)=400grams. I know you plug that into y(t) but how do i isolate for t?

OpenStudy (anonymous):

Plug what into where? You have the diff.eq. \[y'=500-\frac{y}{40+t}\] Have you solve for \(y(t)\) yet?

OpenStudy (anonymous):

i thought you directly plug y(0)=400 into y(t) in 500-(y(t)/(40+t))

OpenStudy (anonymous):

The solution to the diff.eq. is \(y(t)\). First you find a general solution that contains a general constant, then use those initial conditions to solve for the constant.

OpenStudy (anonymous):

o so you find y(t) and then plug those values in? but how do you approach dy/dx=500-(y(t)/(40+t)) when y(t) is apart of it?

OpenStudy (anonymous):

Rearranging the equation a bit, you have \[y'+\frac{1}{40+t}y=500\] Linear first order ODE, so find the integrating factor: \[\large\mu(t)=e^{\int\frac{1}{40+t}~dt}\] Then multiply both sides by \(\mu(t)\).

OpenStudy (anonymous):

.. i have not learned Linear first order... but ill try to multiply both sides by that and see what i get

OpenStudy (anonymous):

Actually, all that stuff I said about the integrating factor isn't necessary. You'll find that \(\mu(t)=40+t\), which (after multiplying both sides of the equation by it) gives you \[(40+t)y'+y=500(40+t)\] Notice that you could have simply multiplied both sides of the equation by the denominator to get the same thing.

OpenStudy (anonymous):

Now the trick is to notice that the left side is \[(40+t)y'+y=\frac{d}{dt}\big[(40+t)y\big]\] So you have \[\frac{d}{dt}\big[(40+t)y\big]=500(40+t)\\ (40+t)y=\int500(40+t)~dt\\ ~~~~~~~~~~~~~~~~~~~\vdots\]

OpenStudy (anonymous):

so (40+t) y = 2000t+250t^2 isolate for y and plug in 400 and solve for t??

OpenStudy (anonymous):

\[(40+t)y=2000t+250t^2+C\\ y(t)=\frac{2000t+250t^2+C}{40+t}\] (That's the general solution above.) Now use the initial condition: y(0)=400: \[400=\frac{2000(0)+250(0)^2+C}{40+(0)}\] Solve for \(C\), then put it back into the general solution.

OpenStudy (anonymous):

oh! i see! thank you so much!!!